# WeTok: Powerful Discrete Tokenization for High-Fidelity Visual Reconstruction

### v2

@@ 4 Experiments
Table 3: Reconstruction evaluation on $\mathbf{256\times 256}$ ImageNet 50K validation set. All models are trained on ImageNet. WeTok achieves SOTA results on different downsampling rates.
@@
ReVQ (Zhang et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $262144$ 2.05 21.96 -
@@
IBQ (Shi et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $262144$ 1.00 20.30 84%
@@
BSQ (Zhao et al., 2024a) 1D $1024$ - $256\times 256$ $2^{18}$ 1.14 25.36 100%
BSQ (Zhao et al., 2024a) 1D 1024 - $256\times 256$ $2^{36}$ 0.45 28.14 100%
@@
Table 4: Zero-shot reconstruction performance on ImageNet 50k validation set and MS-COCO val2017. The tokenizers are trained with large scale general-domain datasets. Text in gray signifies the results directly from Cosmos (Agarwal et al., 2025) report. Our WeTok achieves the best performance on both resolution settings.
@@
Open-MAGVIT2-I-PT (Luo et al., 2024) Discrete 100M 16 341 6.76 22.31 0.65 1.67 22.70 0.64
WeTok (Ours) Discrete 400M 32 768 8.94 20.31 0.55 3.49 20.77 0.55
BSQ (Zhao et al., 2024b) Discrete 1B - 219 - - - 3.81 24.12 0.66
@@
Cosmos (Agarwal et al., 2025) Discrete - 16 384 7.51 20.45 0.52 1.93 20.56 0.51
@@
Open-MAGVIT2-I-PT (Luo et al., 2024) Discrete 100M 16 341 5.10 22.18 0.60 0.78 22.24 0.59
WeTok (Ours) Discrete 400M 32 768 7.83 20.47 0.52 2.03 20.54 0.51
@@
4.1 Implementation
Datasets. To thoroughly evaluate our WeTok, we perform large-scale training on two datasets: (i) 1.2M ImageNet (Russakovsky et al., 2014) training set; (ii) 400M general-domain dataset. To make state-of-the-art comparison, we then evaluate WeTok performance on the ImageNet 50k validation set and MS-COCO 2017 validation set (Lin et al., 2014b). This multi-benchmark evaluation ensures a robust assessment of our model’s capabilities across different domains and data distributions. Unless otherwise stated, we conduct a series of ablation studies on the ImageNet training set. Besides, we train the class-to-image model on the ImageNet training set and test it on the validation set.
Settings. WeTok adopts the architecture proposed in Open-MAGVIT2 (Luo et al., 2024), employing the CNN architecture for encoder, decoder, and discriminator. Images are randomly cropped to $256\times 256$ for training. For the ablation study, all models are trained for 250K steps using a consistent set of hyperparameters. Including a fixed learning rate of 1e-4, the most global batch size is 128, and the Adam (Kingma & Ba, 2014) optimizer with $\beta_{1}=0.5$ and $\beta_{2}=0.9$. For large-scale training, we used the same Adam optimizer settings, but increased the most global batch size to 1024. In this phase, other hyperparameters were individually tuned for each model to achieve optimal performance. For class-to-image generation, we adopt the generative transformer architecture in LlamaGen (Sun et al., 2024). All experiments were carried out using the H20 GPU with Pytorch. More details in Sup. A.
@@ 4.2 Ablation Study
We conducted a comprehensive ablation study to validate the key components of WeTok. We first verified the effectiveness of our two important algorithms, GQ and GD. Then we ablate the performance improvements brought to WeTok by various dimensions, including the number of groups $g$ in GQ, model architecture, training data, and learning rate schedule. In addition, we conduct parameter ablation experiments on the AR model based on WeTok. More implementation details in Sup. A.1. We employ validation loss, rFID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), and PSNR to evaluate the quality of ablation study.
@@
Quantization method. We conduct an ablation study of quantization methods under the same compression ratio, $i.e.$, $d=gd^{\prime}=16$. As shown in Tab. 2, our GQ not only has almost no increase in GPU memory usage due to quantization like BSQ. In Fig. 3, we set the same compression ratio for 3 different quantization methods (LFQ: gg==11, $d^{\prime}$==$16$; BSQ: gg==$16$, $d^{\prime}$==11; GQ: gg==22, $d^{\prime}$==88), GQ performs better than LFQ and far exceeds BSQ.
Figure 10: WeTok-AR-XL generated samples at $\mathbf{256\times 256}$ resolution.
@@
VAR VAR-d16 (Tian et al., 2024) 310M 3.30 274.4 0.84 0.51
VAR-d20 (Tian et al., 2024) 600M 2.57 302.6 0.83 0.56
VAR-d24 (Tian et al., 2024) 1.0B 2.09 312.9 0.82 0.59
VAR-d30 (Tian et al., 2024) 2.0B 1.92 323.1 0.82 0.59
AR LlamaGen-L (Sun et al., 2024) 343M 3.07 256.06 0.83 0.52
@@
Generative decoder. When the performance of the model is saturated after the first stage of reconstruction training, we conduct the second stage of generative training. As shown in Tab. 2, the results show that GD can continue to improve the reconstruction performance of the model, especially in rFID. In addition, we present qualitative ablation results in Fig. 7. After converting the decoder to a generative model, the reconstructed images are more realistic and consistent with visual logic, demonstrating the effectiveness of GD.
Number of group in GQ. We increase the number of groups $g$ by a power of 2. As shown in Fig. 4, the results show that as $g$ increases, the reconstruction performance of the model continues to increase significantly and does not encounter the memory bottleneck like LFQ.
Model architecture. We performed ablation studies to scale up our discrete tokenizer, focusing on the number of base channels and residual blocks in the encoder and decoder. As shown in Fig. 5, across 9 different settings, a configuration with 256 base channels and 4 residual blocks achieves the best reconstruction performance. The encoder and decoder for this optimal architecture contain 198M and 261M parameters, respectively.
Training data. We compare the model trained on the 1.2M ImageNet training set and the other trained on a larger 400M general-domain dataset. As shown in Fig. 6, the model trained on the general-domain dataset achieved a better validation loss and SSIM and PSNR scores. However, it yielded worse rFID and LPIPS scores. We attribute it to the data distribution gap. The ImageNet training and validation sets are in-distribution, whereas the general-domain data are significantly more diverse. This reveals the trade-off between generalization and performance on specific in-distribution evaluation metrics.
Learning rate schedule. A learning rate schedule composed of a warm-up phase followed by a cosine decay is widely adopted in training models. However, we found that this convention may be suboptimal for the training of discrete tokenizers. As illustrated in Fig. 6, the model trained with a constant learning rate demonstrates markedly better performance. Based on this, we exclusively adopted the constant learning rate schedule for all subsequent large-scale training in WeTok.
Parameter of autoregressive model. As shown in Fig. 8, We conduct ablation experiments on the parameter size of the WeTok-based AR model and compared it with an Open-MAGVIT2-based AR model. The experimental results show that the performance of the WeTok-based AR model is slightly inferior to that of the Open-MAGVIT2-based AR model when the parameter size is small. However, as the parameter size increases, the WeTok-based AR model surpasses the Open-MAGVIT2-based AR model in all indicators.
4.3 Comparison with state-of-the-art
Visual Reconstruction. We first evaluate WeTok’s performance in an in-distribution setting, where both the training and validation data originate from the ImageNet dataset. This establishes our model’s core effectiveness before addressing the more complex challenge of distribution gaps. As shown in Tab. 3, WeTok achieves state-of-the-art reconstruction performance, outperforming existing methods across different downsampling ratios. Subsequently, we compare WeTok against other methods, with all models being trained on a large-scale general-domain dataset. To ensure a fair and unbiased comparison, we have carefully filtered this dataset to exclude any data samples in ImageNet and MS-COCO Lin et al. (2014a). This crucial step prevents potential data leakage and ensures that the evaluation accurately measures the models’ ability to generalize from a broad data distribution to the specific target domain. As shown in Tab. 4, our WeTok shows the state-of-the-art reconstruction performance in a wide range of compression ratio scenarios. It not only shows the strongest performance among discrete tokenizers, but also even surpasses the current strongest continuous tokenizers, FLUX-VAE (Labs, 2024) and SD-VAE 3.5 (Esser et al., 2024a). In addition, we present qualitative comparison results on the TokBench (Wu et al., 2025b) in Fig. 9. Compared to the widely used SDXL-VAE (Podell et al., 2023) and SD-VAE 1.5 (Rombach et al., 2022b), WeTok’s reconstruction performance is significantly superior under the same compression ratio.
Visual Generation. To evaluate WeTok’s capabilities beyond image reconstruction, we extend its application to visual generation. We adopt an autoregressive model modified from LlamaGen as Open-MAGVIT2 for this task, primarily due to the well-established scalability of such models. Specifically, we employ the in-distribution WeTok with $16\times$ downsampling in Tab. 3 as the image tokenizer. More training details can be found in Sup. A.2. As shown in Tab. 5, our WeTok-based generative model achieves state-of-the-art performance on the ImageNet 50K validation set. This result demonstrates that our WeTok is a highly effective tokenizer not only for image reconstruction but also for high-fidelity visual generation tasks. As shown in Fig. 10, we show the realistic and diverse image generation results of our WeTok-AR-XL.

### v3

@@ 3.3 Generative Decoder
Figure 8: Parameter ablation experiments of the Autoregressive model with WeTok. The performance of the AR model with WeTok improves more significantly with the increase of parameters.
Table 3: Reconstruction evaluation on $\mathbf{256\times 256}$ ImageNet 50K validation set. All models are trained on ImageNet. WeTok achieves SOTA results on different downsampling rates. * specifies that it is obtained through our testing.
@@
ReVQ (Zhang et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $2^{18}$ 2.05 21.96 -
SweetTok (Yu et al., 2024b) 1D $256$ 16 $256\times 256$ $10481$ 0.73 - -
TiTok (Yu et al., 2024b) 1D $256$ 16 $256\times 256$ $4096$ 1.66 20.01 100%
FlexTok (Bachmann et al., 2025) 1D $256$ - $256\times 256$ $64000$ 1.45 18.53 -
@@
IBQ (Shi et al., 2025) 2D $16\times 16$ 16 $256\times 256$ $2^{18}$ 1.00 20.30 84%
FlowMo-Lo (Sargent et al., 2025) 1D $256$ - $256\times 256$ $2^{18}$ 0.95 22.07 -
@@
BSQ (Zhao et al., 2024b) 1D $1024$ - $256\times 256$ $2^{18}$ 1.14 25.36 100%
FlowMo-Hi (Sargent et al., 2025) 1D $1024$ - $256\times 256$ $2^{18}$ 0.56 24.93 -
Selftok (Wang et al., 2025) 1D $1024$ - $256\times 256$ $2^{15}$ 0.54 26.30 -
BSQ (Zhao et al., 2024b) 1D 1024 - $256\times 256$ $2^{36}$ 0.45 28.14 100%
SweetTok (Yu et al., 2024b) 1D $1024$ 16 $256\times 256$ $10481$ 0.37 - -
@@
Table 4: Zero-shot reconstruction comparison on ImageNet and MS-COCO val2017 validation set. Our WeTok achieves the best performance on both resolution settings.
@@
WeTok (Ours) Discrete 400M 16 384 6.55 21.99 0.63 1.58 22.38 0.62
BSQ (Zhao et al., 2024c) Discrete 1B - 219 - - - 3.81 24.12 0.66
@@
WeTok (Ours) Discrete 400M 32 768 8.94 20.31 0.55 3.49 20.77 0.55
@@
WeTok (Ours) Discrete 400M 16 384 5.30 21.94 0.59 0.81 21.99 0.58
@@
4 Experiments
Datasets. We perform large-scale training on two datasets: (i) 1.2M ImageNet (Russakovsky et al., 2014) training set; (ii) 400M general-domain dataset. For comparison, we evaluate WeTok performance on ImageNet 50k validation set and MS-COCO 2017 validation set (Lin et al., 2014). Unless otherwise stated, we conduct a series of ablation studies on the ImageNet training set. Besides, we train the class-to-image model on the ImageNet training set and test it on the validation set.
Settings. WeTok adopts the architecture proposed in Open-MAGVIT2 (Luo et al., 2024), employing the CNN architecture for encoder, decoder, and discriminator. Images are randomly cropped to $256\times 256$ for training. For ablation study, all models are trained for 250K steps with Adam (Kingma & Ba, 2014) and a consistent set of hyperparameters. For large-scale training, hyperparameters are individually tuned for each model to achieve optimal performance. For class-to-image generation, we adopt the transformer architecture in LlamaGen (Sun et al., 2024). More details in Sup. E.
@@ 4.1 Ablation Study
We conducted a comprehensive ablation study to validate the key components of WeTok. We first verify the effectiveness of our proposed GQ and GD. Then we ablate the performance improvements in various dimensions, including the number of groups, model architecture, training data, and learning rate schedule. In addition, we conduct parameter ablation experiments on the AR model based on WeTok. More details in Sup. E.1. We employ validation loss, rFID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), and PSNR to evaluate the quality of ablation study.
@@
Quantization method. We ablate quantization methods under the same compression ratio, $i.e.$, $dgd^{\prime}$==$16$. As shown in Tab. 2, GQ does not introduce additional memory usage as BSQ. In Fig. 3, we set the same compression ratio for 3 different quantization methods (LFQ: gg==11, $d^{\prime}$==$16$; BSQ: gg==$16$, $d^{\prime}$==11; GQ: gg==22, $d^{\prime}$==88), GQ performs better than LFQ and far exceeds BSQ.
Generative decoder. When the performance of the model is saturated after the first stage of reconstruction training, we conduct the second stage of generative training. As shown in Tab. 2, the results show that GD continues to improve the reconstruction performance of the model, especially in rFID. We present qualitative ablation results in Fig. 7. After converting the decoder to a generative model, the reconstructed images are more realistic, demonstrating the effectiveness of our GD.
Number of groups in GQ. We increase the number of groups $g$ by a power of 2. As shown in Fig. 4, the results show that as $g$ increases, the reconstruction performance of the model continues to increase significantly and does not encounter the memory bottleneck like LFQ.
Model architecture. We ablate the number of base channels and residual blocks in the encoder and decoder to scale up WeTok. As shown in Fig. 5, across 9 different settings, a configuration with 256 base channels and 4 residual blocks achieves the best reconstruction performance. The encoder and decoder for this optimal architecture contain 198M and 261M parameters, respectively.
Training data. As shown in Fig. 6, we ablated models trained on ImageNet versus a large 400M general-domain dataset. The model trained on the general-domain data achieves a lower validation loss and higher SSIM and PSNR, but performs worse on rFID and LPIPS. We attribute to the distribution gap between the general-domain data and the in-distribution ImageNet dataset. This result highlights a trade-off between generalization and performance on in-distribution evaluation metrics.
Learning rate schedule. The warm-up and cosine decay learning rate schedule is widely adopted in training. However, this convention may be suboptimal for the training of discrete tokenizers. In Fig. 6, the model trained with a constant learning rate demonstrates better performance. Based on this, we adopt the constant learning rate schedule for all subsequent large-scale training in WeTok.
Parameter of autoregressive model. As shown in Fig. 8, we ablate the parameter size of the WeTok-based AR model and compared it with Open-MAGVIT2-based AR models. The results show that the performance of the WeTok-based AR model is slightly inferior than the Open-MAGVIT2-based AR model when the parameter size is small. However, as the parameter size increases, WeTok-based AR model surpasses the Open-MAGVIT2-based AR model in all metrics.
4.2 Comparison with state-of-the-art
Visual Reconstruction. We first evaluate WeTok’s performance in in-distribution setting on ImageNet. As shown in Tab. 3, WeTok outperforms existing methods across different downsampling ratios. Subsequently, we evaluate WeTok’s performance in general-domain setting. As shown in Tab. 4, WeTok shows the state-of-the-art reconstruction performance in a wide range of compression ratio scenarios. It not only shows the strongest performance among discrete tokenizers, but also even surpasses the current strongest continuous tokenizers, FLUX-VAE (Labs, 2024) and SD-VAE 3.5 (Esser et al., 2024a). As shown in Fig. 9 and 14, we present qualitative comparison results on the TokBench (Wu et al., 2025b), our WeTok outperforms the widely used SDXL-VAE (Podell et al., 2023) and SD-VAE 1.5 (Rombach et al., 2022b) under the same compression ratio.
Iterative Visual Reconstruction. The use of a tokenizer for iterative image compression-decompression is a promising approach for information transmission (Joshi et al., 2000). As shown in Fig. 10, we compare WeTok (192 compression ratio) with SOTA tokenizers (Labs, 2024; Esser et al., 2024b) on iterative image compression-decompression. We surprisingly find that while reconstructions from leading models like FLUX-VAE and SD-VAE 3.5 collapse after iterations, WeTok’s outputs are remarkably robust and converge to a fixed value. More details in Sup. C.
Visual Generation. To evaluate WeTok’s capabilities for visual generation, we modifys LlamaGen like Open-MAGVIT2. We employ the in-distribution WeTok with $16\times$ downsampling in Tab. 3 as tokenizer. More details in Sup. E.2. As shown in Tab. 5, our WeTok-based AR model achieves state-of-the-art performance on the ImageNet 50K validation set. This result demonstrates that WeTok is a effective tokenizer not only for image reconstruction but also for high-fidelity visual generation. As shown in Fig. 15, we show the realistic and diverse image generation results of our WeTok-AR-XL.
Figure 10: Qualitative and quantitative comparison of iterative image reconstruction.
@@ 4.2 Comparison with state-of-the-art
LlamaGen-L (Sun et al., 2024) 343M 3.07 256.06 0.83 0.52